OpenStudy (anonymous):

check my work pls lim x=> 4 from the left abs(x-4)/x-4

OpenStudy (anonymous):

is the answer just 1?

OpenStudy (anonymous):

No, I don't think so. The function is not continuous from the left side

OpenStudy (anonymous):

how would you solve it then?

OpenStudy (freckles):

Consider the left and right limit; are they the same. Hint abs( x-4) =x-4 if x-4 >0 and abs( x-4) =-(x-4) if x-4<0

OpenStudy (freckles):

Oh I didn't see the left lol Though you were looking for actual Just consider the what happens to the left that is you are looking at x<4

OpenStudy (freckles):

Use what I said before about the and value part

OpenStudy (freckles):

Abs*

OpenStudy (anonymous):

abs(x-4)=x-4 if x-4>0?

OpenStudy (freckles):

We don't want to use that That is for values to the right of 4. We are looking to the left so days your question. We want to use the abs(x - 4)=-(x-4) since x<4

OpenStudy (anonymous):

what I did was as x<4, abs(x-4) = x-4 therefore lim =>4- abs(x-4)/x-4 so lim x=> 4- x-4/x-4 = 1

OpenStudy (freckles):

You are actually finding the right limit because you used abs(x-4)=x-4 for when x>4

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

so it would be -1?

OpenStudy (freckles):

Again abs(x-4)=-(x-4) since x<4 and yes - (x-4)/(x-4)=1 when x goes to four from legy

OpenStudy (freckles):

-1

OpenStudy (freckles):

Sorry on phone hard to typet

OpenStudy (anonymous):

hehe thanks

OpenStudy (freckles):

Lol but I think you get this now just remember to consider the abs value function as a piece wise function

OpenStudy (anonymous):

can you give me another examples pls?

OpenStudy (anonymous):

*example

OpenStudy (freckles):

Abs(f(x)) =f(x) if f(x) >0 and abs(f(x)) =-f(x) if f(x) <0

OpenStudy (anonymous):

so in other words you are saying that the answer would be = +-1 if they were asking for lim x=>4 ?

OpenStudy (freckles):

Right=1 Left=-1 Actusl doesn't exist because left doesn't equal right

OpenStudy (anonymous):

hmm

OpenStudy (freckles):

For right you would have (x-4)/(x-4)=1 and left we had - (x-4)/(x-4)=-1

OpenStudy (anonymous):

yep i think i got it now

OpenStudy (freckles):

If I was on my desktop I would give you another example but this phone doesn't make answering or replying to questions easy

OpenStudy (anonymous):

you can send me another example when you get to your desktop! :D thanks for your help

OpenStudy (freckles):

No problem